Session: Feb 1998 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (tony foley at Sat Feb 14 20:58:30 MET 1998) Minor failures
Before start the vacuum motor blew a fuse, but we found a replacement and got the tape working about 20UT

Medicina (Morsiani at Sun Feb 15 10:07:47 MET 1998) Success

Jodrell (Alastair Gunn at Mon Feb 16 11:13:47 MET 1998) Minor failures
FS reporting tape head 2 in wrong position for some scans - investigating. No other known problems.

Onsala (John Conway at Mon Feb 16 12:00:47 MET 1998) Success
No known problems. However we used our new Septum polariser for this experiment. We mounted this for the previous dual polarisation cal expt as a test,and intended to switch back afterward to our old quarter wave plate polariser but weather prevented safe access to focus box. We made our best effort to ensure that the polariser was correctly orientated but this is the first time used in a VLBI expt...

Effelsberg (dgraham at Mon Feb 16 15:51:39 MET 1998) Minor failures
sources too low elevation after 0810 UT.

Torun (K.Borkowski at Mon Feb 16 16:28:05 MET 1998) Success
No known problems

Noto (C.Trigilio at Fri Feb 20 15:39:53 MET 1998) Success

Calibration (jean\-francois Desmurs at Mon Mar 2 18:45:12 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb98 directory under the name gg034.antab. Read the comments in the file to get further information.